It usually takes Eva 3 hours longer to do the monthly payroll than it takes Cindy. They start working on it together at 9.00 AM and at 5.00 PM they have 90% of it done. If Eva took s 2 hour lunch break while Cindy had none, then how much longer will it take for them to finish the payroll working together?''.
Found 2 solutions by DRichardson, stanbon:
Answer by DRichardson(63)

It usually takes Eva 3 hours longer to do the monthly payroll than it takes Cindy. They start working on it together at 9.00 AM and at 5.00 PM they have 90% of it done. If Eva took s 2 hour lunch break while Cindy had none, then how much longer will it take for them to finish the payroll working together?''

Hello there,
What you would do here is you would find how much they got done in each hour, and then take two hours away... let me show you the answer:
10% per hour
20% would be gone if Eva took a two hour lunch break
So when eva takes that lunch break, 20% of the payroll time is lost.
70% of the work is done in 9 hours
It woudl take them three more hours to finish it because if they go at a rate of 10% per hour , and they have 30 % to go, then three hours is the answer
Does this help?
Delilah R.

Answer by stanbon(75887)

It usually takes Eva 3 hours longer to do the monthly payroll than it takes Cindy. They start working on it together at 9.00 AM and at 5.00 PM they have 90% of it done. If Eva took s 2 hour lunch break while Cindy had none, then how much longer will it take for them to finish the payroll working together?
--------------------------------------
Cindy's rate is 1/x job/hr
-------------
Eva's rate is 1/(x+3) job/hr
-------------------------------
From 9am to 5pm they have worked 8 hrs
In that time they have accomplished
8*Cindy's rate + 6*Eva's rate = 90% of the job
8(1/x)+6 = 0.90 job
I get x = 24.7
-------------------------------------
Solve that for "x"; then you will know Cindy
and Eva's rates which are 1/x = 1/24.7 and 1/27.7).
---------------------
To get the time remaining for them to finish
the job.
k(1/24.7 + 1/27.7) = 0.10 of the job
k(0.0766) = 0.10
k = 1.306 hrs.
============================
Cheers,
Stan H.